Operations Report – January 22nd

Crew 203 Operations Report 22-JAN-2019

SOL: 10

Name of person filing report: Freddy Castaneda

Non-nominal systems: Nothing new to report.

Notes on non-nominal systems: None

Generator (hours run): 9hr 30min; Turned on last night (21Jan2019) at 17:40; Stopped this morning (22Jan2019) at 3:10; Turned on tonight (22Jan2019) at 16:50

Solar SOC – Turned on (21Jan2019) 86%; Turned off (22Jan2019) 100%; Turned on (22Jan2019) 83%

Diesel Reading – Under 1/2 tank.

Propane Reading – About 19%

Ethanol Free Gasoline – Nothing to report.

Water (auxiliary tank) – Not in use.

Water (static tank) – About 35%; 190 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not applicable

Water in GreenHab – About 49%; 147 gallons

Water (loft) – At level marker 12

Static to Loft Pump used – Yes; At 18:40 to refill tank

Water Meter: 01403746

Toilet tank emptied: No.

Deimos rover used: No, still not functional

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Used for EVA to Pooh´s Corner.

Hours: 69.7

Beginning charge: 100%

Ending charge: 86%

Currently charging: Yes

Opportunity rover used: Used for EVA to Pooh´s Corner.

Hours: 46.2

Beginning charge: 100%

Ending charge: 84%

Currently charging: Yes

Curiosity rover used: Used for EVA to Pooh´s Corner.

Hours: 71.6

Beginning charge: 100%

Ending charge: 95%

Currently charging: Yes

Notes on rovers: Brakes on Opportunity still not working properly.

ATV’s Used: None (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: None

Oil Added? No

ATV Fuel Used: None

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? Not used.

CrewCar used and why, where? Off campus

General notes and comments: Nothing to report.

Summary of internet: Nothing to report.

Summary of suits and radios: Suit #3 still pending for review.

Summary of Hab operations: Nothing to report.

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: Nothing to report.